# Stability-Certified On-Policy Data-Driven LQR via Recursive Learning and Policy Gradient

### v1

@@ 4.2 Averaged System Analysis
Next, we carry out the stability analysis of the time-varying system (4.1) by leveraging on the averaging and singular perturbations theories (cf. Appendix A for further details). Indeed, since system (4.1) enjoys a two-time-scale structure (cf. the generic system (A) in Appendix A), we can study (4.1) by only investigating an auxiliary system typically termed as the averaged system. The latter is obtained by considering the slow dynamics (32b) in which (i) the fast state is frozen to its equilibrium, i.e., with $\tilde{\chi}_{t}=0$ for all $t\geq 0$, and (ii) the vector field describing the dynamics is averaged with respect to time.
@@
Let the assumptions of Theorem 3.2 hold true. Consider $f$ defined in (33j). Then, it holds
$\lim_{T\to\infty}\frac{1}{T}\sum_{\tau=\bar{t}+1}^{\bar{t}+T}f(0,\tilde{\mathrm{z}},\tau)=\!-\!\begin{bmatrix}\tilde{\theta}\\ G(\tilde{K}\!+\!K^{\star},\tilde{\theta}\!+\!\theta^{\star})\end{bmatrix}$ (39)
uniformly in $\bar{t}\in\mathbb{N}$ and for all $\tilde{\mathrm{z}}=\operatorname{col}(\tilde{\theta},\tilde{K})\in\mathbb{R}^{n_{\mathrm{z}}}$. normal-$\blacksquare$
The proof of Lemma 4.2 is given in Appendix C.
Lemma 4.2 provides a suitable approximation of the dynamics of $\tilde{\mathrm{z}}$ in (32b) when (i) the convergence of the fast state $\tilde{\chi}$ to its equilibrium has already occurred and (ii) by averaging over time $t$ the vector field $f(0,\tilde{\mathrm{z}},t)$. Specifically, under this approximation, Lemma 4.2 ensures that the two components of the driving term of the dynamics of $\tilde{\mathrm{z}}$ are given by (i) a proportional term $-\gamma\tilde{\theta}$ and (ii) an approximate version of the correct gradient $G(\tilde{K}+K^{\star},\theta^{\star})$. Next, we will leverage averaging theory to prove the stability of the origin for system (4.1).
@@
$f^{\textsc{av}}(\tilde{\mathrm{z}})\!:=\!\lim_{T\to\infty}\!\frac{1}{T}\!\sum_{\tau=\bar{t}+1}^{\bar{t}+T}\!\!f(0,\tilde{\mathrm{z}},t)\!=\!-\begin{bmatrix}\tilde{\theta}\\ G(\tilde{K}+K^{\star},\tilde{\theta}+\theta^{\star})\end{bmatrix}\!\!,$
in which $\tilde{\mathrm{z}}=\operatorname{col}(\tilde{\theta},\tilde{K})$. Then, we define the averaged system, with state $\tilde{\mathrm{z}}^{\textsc{av}}_{t}\in\mathbb{R}^{n_{\mathrm{z}}}$, associated to (4.1) as
@@
Exploding the expression of $f^{\textsc{av}}$ (cf. (39)) and $\tilde{\mathrm{z}}_{t}:=\operatorname{col}(\tilde{\theta}_{t},\tilde{K}_{t})$, the dynamics in (40) results in a cascade as depicted in Figure 4.
@@
Let the assumptions of Theorem 3.2 hold true. Consider the averaged system (40). Then, for all $\tilde{\mathrm{z}}^{\textsc{av}}_{0}\in\mathbb{R}^{(n+2m)\times n}$ such that the corresponding $A^{\textsc{av}}_{0}+B^{\textsc{av}}_{0}(\tilde{K}^{\textsc{av}}_{0}+K^{\star})$ and $A+B(\tilde{K}^{\textsc{av}}_{0}+K^{\star})$ are Schur matrices, there exists $\bar{\gamma}^{\textsc{av}}>0$ such that, for all $\gamma\in(0,\bar{\gamma}^{\textsc{av}})$, the origin of (40) is exponentially stable. normal-$\blacksquare$
The proof of Proposition 4.3 is given in Appendix D.
@@
We will use Theorem A.6 given in Appendix A to guarantee the exponential stability of the origin for (4.1). Specifically, in order to apply Theorem A.6, we need to verify
@@
    the exponential stability of the origin for the associated averaged system,
@@
    the Lipschitz continuity of the vector field of (4.1),
@@
    that the origin is an equilibrium point of (4.1), and
@@
    that
    $\left\|\dfrac{1}{T}\sum_{\tau=\bar{t}+1}^{\bar{t}+T}\Delta f(\tilde{\mathrm{z}},\tau)\right\|\leq\nu(T)\left\|\tilde{\mathrm{z}}\right\|$ $\left\|\dfrac{1}{T}\sum_{\tau=\bar{t}+1}^{\bar{t}+T}\frac{\partial\tilde{f}(\tilde{\mathrm{z}},\tau)}{\partial\tilde{\mathrm{z}}}\right\|\leq\nu(T),$ (43)
    where $\Delta f(\tilde{\mathrm{z}},\tau):=f(0,\tilde{\mathrm{z}},\tau)-f^{\textsc{av}}(\tilde{\mathrm{z}})$ and $\nu(t)$ is a nonnegative strictly decreasing function with the property $\nu(t)\to 0$ as $t\to\infty$.
As for condition (i), it follows from Proposition 4.3. Condition (ii) is satisfied by using the quantities defined in (D) in Appendix D as the required Lipschitz constants of the vector field of (4.1). Condition (iii) can be verified by means of (36). Finally, in order to check condition (iv) (cf. (43)), note that
$\tilde{f}(\tilde{\mathrm{z}},t)=(H^{{\textnormal{ss}}}_{t})^{\dagger}H^{{\textnormal{ss}}}_{t}\tilde{\theta}-\tilde{\theta}\stackrel{{\scriptstyle(a)}}{{=}}0,$ (44)
where in $(a)$ we use the fact that $H^{{\textnormal{ss}}}_{t}$ is invertible for all $t\in N$ (cf. Lemma C.1 in Appendix D). Therefore, the conditions in (43) are satisfied and, thus, we can apply Theorem A.6. This result guarantees the existence of $\bar{\gamma}>0$ such that, for all $\gamma\in(0,\bar{\gamma})$, the origin is an exponentially stable equilibrium point for system (4.1). The proof follows backtracking to the original coordinates $(\mathrm{x},\theta,K)$.
@@ 5 Numerical Simulations
$\begin{split}\dot{x}=&\left[\begin{smallmatrix}-0.0151&-60.5651&0&-32.174\\ -0.0001&-1.3411&0.9929&0\\ 0.00018&43.2541&-0.86939&0\\ 0&0&1&0\end{smallmatrix}\right]x+\left[\begin{smallmatrix}-2.516&-13.136\\ -0.1689&-0.2514\\ -17.251&-1.5766\\ 0&0\end{smallmatrix}\right]u,\end{split}$ (45)
where the state $x\in\mathbb{R}^{4}$ represents the forward velocity, the attack angle, the pitch rate and the pitch angle, while the inputs $u\in\mathbb{R}^{2}$ are the elevator and flaperon angles. The discrete-time system matrices ${A}_{\star}$ and ${B}_{\star}$ of (1d) are discretized from the continuous-time system (45) using a Zero Order Hold on the input with sampling time $T_{s}=0.05$ [[[[s]]]]. Notice that the resulting matrix ${A}_{\star}$ has one eigenvalue outside the unit disk, i.e., it is not Schur. The cost matrices $Q\in\mathbb{R}^{4\times 4}$ and $R\in\mathbb{R}^{2\times 2}$ are randomly generated, while ensuring that $Q=Q^{\top}\geq 0$ and $R=R^{\top}>0$. We set $\gamma=10^{-4}$ and $\left\|\mathrm{w}_{0}\right\|=0.01$.
5.1 Exogenus System Design Procedure
Before providing the results of the numerical simulations, we propose a procedure tailored to design an exogenous system such that Assumption 3.1 is guaranteed. For the sake of completeness, we consider the general case with the state dimension being ${n}$ and the input dimension being ${m}$. First, we set $q:=n+1$, $n_{\mathrm{w}}=qm$, and $F:=\operatorname{blkdiag}(F_{1},\dots,F_{q}),$ where, for all $i\in\{1,\ldots,q\}$,
for some given $\omega_{i}>0$ such that $\omega_{i}$ and $\omega_{j}$ are uncorrelated for each pair $i,j\in\{1,\ldots,q\}$ with $i\neq j$. By choosing an initial condition $\mathrm{w}_{0}$ that satisfies
$([\mathrm{w}_{0}]_{2i+1})^{2}+([\mathrm{w}_{0}]_{2i+2})^{2}\neq 0,\quad 0\leq i\leq q-1,$ (46)
the chosen structure of $F$ guarantees (18a) according to [53, Th.2]. As for (18f), we achieve it by selecting $E$ such that $\begin{bmatrix}E^{\top}&(EF)^{\top}&\ldots&(EF^{n})^{\top}\end{bmatrix}$ is nonsingular.
@@ 5.2 Aircraft Control
We start by considering the LTI system (45). We run Relearn LQR with the exogenus signal generated via the procedure detailed above. In Figure 6 (left) it is possible to observe the evolution of the normalized cost error $|J(K_{t},\theta^{\star}_{t})-J^{\star}|/J^{\star}$, with $J^{\star}:=J(K^{\star},\theta^{\star})$ and $\theta^{\star}:=[{A}_{\star}\;{B}_{\star}]^{\top}$, in logarithmic scale. On the right of Figure 6 it is depicted the evolution of the normalized estimation error $\left\|\theta_{t}-\theta^{\star}\right\|/\left\|\theta^{\star}\right\|$ in logarithmic scale. Notice that, in both cases, convergence to the optimal cost $J^{\star}$ and true parameters $\theta^{\star}$ is achieved. Finally, in Figure 5 the state trajectory of the closed-loop system is depicted. The initial condition $\mathrm{x}_{0}$ is sampled from a normal distribution with mean value $10$ for each state. Notice that, after a transient, the states oscillate about the origin due to the exogenous system.
@@ 5.3 Aircraft Control with Drifting Parameters
Figure 8: (Left) Evolution of the normalized cost error $|J(K_{t},\theta^{\star}_{t})-J^{\star}_{t}|/J^{\star}_{t}$. (right) Evolution of the normalized estimation error $\left\|\theta_{t}-\theta^{\star}_{t}\right\|/\left\|\theta^{\star}_{t}\right\|$ (left).
@@ 6 Conclusions
Appendix A Preliminaries on averaging theory for two-time-scale systems
We report [54, Theorem 2.2.4], which is a useful result in the context of averaging theory for two-time-scale systems. Consider the time-varying system
$\chi_{t+1}$ $=\mathcal{A}(\mathrm{z}_{t})\chi_{t}+h(\mathrm{z}_{t},t)+\epsilon g(\chi_{t},\mathrm{z}_{t},t)$ (A.1a)
$\mathrm{z}_{{t+1}}$ $=\mathrm{z}_{t}+\epsilon f(\chi_{t},\mathrm{z}_{t},t),$ (A.1b)
with $\chi_{t}\in\mathbb{R}^{n_{\chi}}$, $\mathrm{z}_{t}\in\mathbb{R}^{n_{\mathrm{z}}}$, $g:\mathbb{R}^{n_{\chi}}\times\mathbb{R}^{n_{\mathrm{z}}}\times\mathbb{N}\to\mathbb{R}^{n_{\chi}}$, $f:\mathbb{R}^{n_{\chi}}\times\mathbb{R}^{n_{\mathrm{z}}}\times\mathbb{N}\to\mathbb{R}^{n_{\chi}}$, and $A:\mathbb{R}^{n_{\mathrm{z}}}\to\mathbb{R}^{n_{\chi}\times n}$. We enforce the following assumptions.
Assumption A.1
There exists $r$ such that $f$, $g$, and $h$ are Lipschitz continuous into $\mathcal{B}_{r}(0_{n_{\chi}+n_{\mathrm{z}}})$. normal-$\blacksquare$
Assumption A.2
It holds $h(0,t)=0$, $g(0,0,t)=0$, and $f(0,0,t)=0$ for all $t\in\mathbb{N}$. normal-$\blacksquare$
Assumption A.3
Moreover, there exists $k_{a}>0$ such that
for all $i\in\{1,\ldots,m\}$ and $\mathrm{z}\in\mathcal{B}_{r}(0_{n_{\mathrm{z}}})$. normal-$\blacksquare$
Assumption A.4
The function $f$ is piecewise continuous in $t$ with the limit
$f^{\textsc{av}}(\mathrm{z}):=\lim_{T\to\infty}\frac{1}{T}\sum_{\tau=\bar{t}+1}^{\bar{t}+T}f(0,\mathrm{z},\tau)$ (A.2)
existing uniformly in $\bar{t}\in\mathbb{N}$ and for all $\mathrm{z}\in\mathcal{B}_{r}(0_{n_{\mathrm{z}}})$. normal-$\blacksquare$
We associate a so-called averaged system to (A) given by
$\mathrm{z}^{\textsc{av}}_{t+1}=\mathrm{z}^{\textsc{av}}_{t}+\epsilon f^{\textsc{av}}(\mathrm{z}^{\textsc{av}}_{t}).$ (A.3)
Assumption A.5
Consider $f^{\textsc{av}}$ as defined in (A.2) and let $\Delta f:\mathbb{R}^{n_{\mathrm{z}}}\times\mathbb{N}\to\mathbb{R}^{n_{\mathrm{z}}}$ be defined as
$\Delta f(\mathrm{z},t):=f(0,\mathrm{z},t)-f^{\textsc{av}}(\mathrm{z}).$
$\left\|\dfrac{1}{T}\sum_{\tau=\bar{t}+1}^{\bar{t}+T}\tilde{f}(\mathrm{z},\tau)\right\|$ $\leq\nu(T)\left\|\mathrm{z}\right\|$
$\left\|\dfrac{1}{T}\sum_{\tau=\bar{t}+1}^{\bar{t}+T}\frac{\partial\tilde{f}(\mathrm{z},\tau)}{\partial\mathrm{z}}\right\|$ $\leq\nu(T),$
uniformly in $\bar{t}\in\mathbb{N}$ and for all $\mathrm{z}\in\mathcal{B}_{r}(0_{n_{\mathrm{z}}})$. normal-$\blacksquare$
Theorem A.6
[54, Theorem 2.2.4] Consider system (A) and let Assumptions A.1A.2A.3A.4 and A.5 hold. If there exists $\epsilon_{0}>0$ such that, for all $\epsilon\in(0,\epsilon_{0})$, the origin is exponentially stable for system (A.3), then there exists $\epsilon_{1}>0$ such that, for all $\epsilon\in(0,\epsilon_{1})$, the origin is an exponentially stable equilibrium of system (A). normal-$\blacksquare$

### v2

@@ 4.2 Averaged System Analysis
Next, we carry out the stability analysis of the time-varying system (30) by leveraging on the averaging and timescale separation theories (cf. Section 2.1 for further details). Indeed, since system (30) enjoys a two-time-scale structure (cf. the generic system (1) in Section 2.1), we can study (30) by only investigating an auxiliary system typically termed as the averaged system. The latter is obtained by considering the slow dynamics (30b) in which (i) the fast state is frozen to its equilibrium, i.e., with $\tilde{\chi}_{t}=0$ for all $t\geq 0$, and (ii) the vector field describing the dynamics is averaged with respect to time.
@@
Let the assumptions of Theorem 3.3 hold true. Consider $f$ defined in (31c). Then, it holds
$\lim_{T\to\infty}\frac{1}{T}\sum_{\tau=\bar{t}+1}^{\bar{t}+T}f(0,\tilde{\mathrm{z}},\tau)=\!-\!{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\begin{bmatrix}G(\tilde{K}\!+\!K^{\star},\tilde{\theta}\!+\!\theta^{\star})\\ \tilde{\theta}\end{bmatrix}}$ (35)
uniformly in $\bar{t}\in\mathbb{N}$ and for all $\tilde{\mathrm{z}}=\operatorname{col}(\tilde{\theta},\tilde{K})\in\mathbb{R}^{n_{\mathrm{z}}}$. $\blacksquare$
The proof of Lemma 4.2 is given in Appendix B.
Lemma 4.2 provides a suitable approximation of the dynamics of $\tilde{\mathrm{z}}$ in (30b) when (i) the convergence of the fast state $\tilde{\chi}$ to its equilibrium has already occurred and (ii) by averaging over time $t$ the vector field $f(0,\tilde{\mathrm{z}},t)$. Specifically, under this approximation, Lemma 4.2 ensures that the two components of the driving term of the dynamics of $\tilde{\mathrm{z}}$ are given by (i) a proportional term $-\gamma\tilde{\theta}$ and (ii) an approximate version of the correct gradient $G(\tilde{K}+K^{\star},\theta^{\star})$. Next, we will leverage averaging theory to prove the stability of the origin for system (30).
@@
$f^{\textsc{av}}(\tilde{\mathrm{z}})\!:=\!\lim_{T\to\infty}\!\frac{1}{T}\!\sum_{\tau=\bar{t}+1}^{\bar{t}+T}\!\!f(0,\tilde{\mathrm{z}},t).$
Then, we define the averaged system, with state $\tilde{\mathrm{z}}^{\textsc{av}}_{t}\in\mathbb{R}^{n_{\mathrm{z}}}$, associated to (30) as
@@
Exploding the expression of $f^{\textsc{av}}$ (cf. (35)) and $\tilde{\mathrm{z}}^{\textsc{av}}_{t}:=\operatorname{col}(\tilde{\theta}^{\textsc{av}}_{t},\tilde{K}^{\textsc{av}}_{t})$, the dynamics in (36) results in a cascade as depicted in Figure 4. Such a cascade structure motivates the unnecessity of a timescale separation between the dynamics of $K_{t}$ and $\theta_{t}$, therefore their updates (cf. (12c) and (13)) share a common parameter $\gamma$.
@@
Let the assumptions of Theorem 3.3 hold true. Consider the averaged system (36). Then, for all $\tilde{\mathrm{z}}^{\textsc{av}}_{0}\in\mathbb{R}^{(n+2m)\times n}$ such that the corresponding $A^{\textsc{av}}_{0}+B^{\textsc{av}}_{0}(\tilde{K}^{\textsc{av}}_{0}+K^{\star})$ and $A+B(\tilde{K}^{\textsc{av}}_{0}+K^{\star})$ are Schur matrices, there exists $\bar{\gamma}^{\textsc{av}}>0$ such that, for all $\gamma\in(0,\bar{\gamma}^{\textsc{av}})$, the origin of (36) is exponentially stable. $\blacksquare$
The proof of Proposition 4.3 is given in Appendix C.
@@
We will use Theorem 2.6 given in Section 2.1 to guarantee the exponential stability of the origin for (30). Specifically, in order to apply Theorem 2.6, we need to verify
@@
    the exponential stability of the origin for the averaged system (36);
@@
    the Lipschitz continuity of the vector field of the original system (30);
@@
    that the origin is an equilibrium point of the original system (30);
@@
    that the matrix function $\mathcal{A}(\tilde{\mathrm{z}})$ is uniformly Schur in a ball $\mathcal{B}_{r}(0_{n_{\mathrm{z}}})$;
  5. (v)
    the following two conditions for the difference between the vector field $f$ of the original system (30b) and the vector field $f^{\textsc{av}}$ of the reduced system (36), namely
    $\left\|\dfrac{1}{T}\sum_{\tau=\bar{t}+1}^{\bar{t}+T}\Delta f(\tilde{\mathrm{z}},\tau)\right\|\leq\nu(T)\left\|\tilde{\mathrm{z}}\right\|$ (38a)
    $\left\|\dfrac{1}{T}\sum_{\tau=\bar{t}+1}^{\bar{t}+T}\frac{\partial{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Delta f}(\tilde{\mathrm{z}},\tau)}{\partial\tilde{\mathrm{z}}}\right\|\leq\nu(T),$ (38b)
    for all $\tau\in\mathbb{N}$, where $\Delta f(\tilde{\mathrm{z}},\tau):=f(0,\tilde{\mathrm{z}},\tau)-f^{\textsc{av}}(\tilde{\mathrm{z}})$ and $\nu(t)$ is a nonnegative strictly decreasing function with the property $\nu(t)\to 0$ as $t\to\infty$.
As for condition (i), it follows from Proposition 4.3. Condition (ii) is satisfied by using the quantities defined in (C.3) in Appendix C as the required Lipschitz constants of the vector field of (30). Condition (iii) can be verified by means of (34). As for condition (iv), we note that the matrix $\mathcal{A}(\tilde{\mathrm{z}})$ is Schur for all $\tilde{K}\in\mathbb{R}^{m\times n}$ (the first component of $\tilde{\mathrm{z}}$, see its definition in (29)) such that $(\tilde{K}+K^{\star})\in\mathcal{K}$ by definition of $\mathcal{K}$ as the set of stabilizing gains for the pair $({A}_{\star},{B}_{\star})$ (see (8)). Hence, condition (iv) is verified with the largest ball contained in $\mathcal{K}$ and centered in $K^{\star}$. Finally, in order to check condition (v) (cf. (38)), note that
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Delta f}(\tilde{\mathrm{z}},t)=(H^{{\textnormal{ss}}}_{t})^{\dagger}H^{{\textnormal{ss}}}_{t}\tilde{\theta}-\tilde{\theta}=0,$ (39)
where in we used the fact that $H^{{\textnormal{ss}}}_{t}$ is invertible for all $t\in N$ (cf. Lemma B.1 in Appendix C). Therefore, the conditions in (38) are satisfied and, thus, we can apply Theorem 2.6. This result guarantees the existence of $\bar{\gamma}>0$ such that, for all $\gamma\in(0,\bar{\gamma})$, the origin is an exponentially stable equilibrium point for system (30). The proof follows backtracking to the original coordinates $(\mathrm{x},H,S,\theta,K)$.
@@ 5 Numerical Simulations
$\begin{split}\dot{x}=&\left[\begin{smallmatrix}-0.0151&-60.5651&0&-32.174\\ -0.0001&-1.3411&0.9929&0\\ 0.00018&43.2541&-0.86939&0\\ 0&0&1&0\end{smallmatrix}\right]x\\ &\hskip 56.9055pt+\left[\begin{smallmatrix}-2.516&-13.136\\ -0.1689&-0.2514\\ -17.251&-1.5766\\ 0&0\end{smallmatrix}\right]u,\end{split}$ (40)
where the state $x\in\mathbb{R}^{4}$ represents the forward velocity, the attack angle, the pitch rate and the pitch angle, while the inputs $u\in\mathbb{R}^{2}$ are the elevator and flaperon angles. The discrete-time system matrices ${A}_{\star}$ and ${B}_{\star}$ of (5b) are discretized from the continuous-time system (40) using a Zero Order Hold on the input with sampling time $T_{s}=0.05$ [[[[s]]]]. Notice that the resulting matrix ${A}_{\star}$ has one eigenvalue outside the unit disk, i.e., it is not Schur. The cost matrices $Q\in\mathbb{R}^{4\times 4}$ and $R\in\mathbb{R}^{2\times 2}$ are randomly generated, while ensuring that $Q=Q^{\top}\geq 0$ and $R=R^{\top}>0$. As for the design of the matrices $F$ and $E$, we used the procedure outlined in Remark 3.2 by choosing $\omega_{1}=1/5$, $\omega_{i+1}=\omega_{i}$ if $i$ is odd and $\omega_{i}=2\omega_{i-2}$ if $i$ is even for all $i\in\{2,\dots,n_{\mathrm{w}}\}$. Further, we set $\gamma=5\cdot 10^{-4}$ and $\mathrm{w}_{0}=\begin{bmatrix}0.01&0&\dots&0.01&0\end{bmatrix}^{\top}$.
@@ 5.1 Aircraft Control
We start by considering the LTI system (40). We run Relearn LQR with the exogenous signal generated via the procedure detailed above. In Figure 6 (left) it is possible to observe the evolution of the normalized cost error $|J(K_{t},\theta^{\star}_{t})-J^{\star}|/J^{\star}$, with $J^{\star}:=J(K^{\star},\theta^{\star})$ and $\theta^{\star}:=[{A}_{\star}\;{B}_{\star}]^{\top}$, in logarithmic scale. On the right of Figure 6 it is depicted the evolution of the normalized estimation error $\left\|\theta_{t}-\theta^{\star}\right\|/\left\|\theta^{\star}\right\|$ in logarithmic scale. Notice that, in both cases, convergence to the optimal cost $J^{\star}$ and true parameters $\theta^{\star}$ is achieved. Finally, in Figure 5 the state trajectory of the closed-loop system is depicted. The initial condition $\mathrm{x}_{0}$ is sampled from a normal distribution with mean value $10$ for each state. Notice that, after a transient, the states oscillate about the origin due to the exogenous system.
@@ 5.2 Aircraft Control with Drifting Parameters
Figure 8: Evolution of the normalized cost error $|J(K_{t},\theta^{\star}_{t})-J^{\star}_{t}|/J^{\star}_{t}$ (left). Evolution of the normalized estimation error $\left\|\theta_{t}-\theta^{\star}_{t}\right\|/\left\|\theta^{\star}_{t}\right\|$ (right).
@@ 6 Conclusions
Although our strategy is analyzed with a time-invariant plant without disturbances, the procedure we developed is inherently applicable to more challenging scenarios where, for instance, the system and cost matrices vary over time and the plant is affected by disturbances. Thus, our work lays the foundation for obtaining stability certificates also in such complex settings.
Appendix A Proof of Lemma 4.1
We note that (27) is obtained by setting $K_{t}=K^{\star}$ in (24) (which compactly collects the updates (19a), (19b), and (23)). Hence, we start by inspecting (19a) and (19b) restricted to the manifold in which $K_{t}=K^{\star}$, namely
$\mathrm{w}_{{t+1}}$ $=F\mathrm{w}_{t}$ (A.1a)
$\mathrm{\mathrm{x}}_{{t+1}}$ $=({A}_{\star}+{B}_{\star}K^{\star})\mathrm{\mathrm{x}}_{t}+{B}_{\star}E\mathrm{w}_{t}.$ (A.1b)
System (A.1) is a cascade, therefore its steady-state solution is $\operatorname{col}(\mathrm{w}_{t},\mathrm{\mathrm{x}}_{t})=\operatorname{col}(I_{n_{\mathrm{w}}},\Pi_{\mathrm{x}})\mathrm{w}_{t}$, with $\Pi_{\mathrm{x}}\in\mathbb{R}^{n\times m}$ solution to the following Sylvester equation
$\Pi_{\mathrm{x}}F=({A}_{\star}+{B}_{\star}K^{\star})\Pi_{\mathrm{x}}+{B}_{\star}E.$ (A.2)
Being $F$ marginally stable (cf. Assumption 3.1) and ${A}_{\star}+{B}_{\star}K^{\star}$ Schur (so that $\sigma(F)\cap\sigma({A}_{\star}+{B}_{\star}K^{\star})=\emptyset$) the solution $\Pi_{\mathrm{x}}$ exists and is unique. Then, we inspect the dynamics (23) restricted to the manifold in which $\mathrm{\mathrm{x}}_{t}=\Pi_{\mathrm{x}}\mathrm{w}_{t}$ and $K_{t}=K^{\star}$. Let $M\in\mathbb{R}^{(n+m)\times(n+m)}$ be
$M:=\begin{bmatrix}\Pi_{\mathrm{x}}^{\top}&(K^{\star}\Pi_{\mathrm{x}}+E)^{\top}\end{bmatrix}^{\top},$ (A.3)
then it holds
$\mathtt{vec}\left(\mathrm{w}_{{t+1}}\mathrm{w}_{{t+1}}^{\top}\right)=\mathtt{vec}\left(F\mathrm{w}_{t}\mathrm{w}_{t}^{\top}F^{\top}\right)$ (A.4a)
$H^{\text{vc}}_{t+1}=\lambda H^{\text{vc}}_{t}+\mathtt{vec}\left(M\mathrm{w}_{t}\mathrm{w}_{t}^{\top}M^{\top}\right)$ (A.4b)
$S^{\text{vc}}_{t+1}=\lambda S^{\text{vc}}_{t}+\mathtt{vec}\left(M\mathrm{w}_{t}\mathrm{w}_{t}^{\top}M^{\top}\theta^{\star}\right),$ (A.4c)
where the first equation comes from the vectorization of (19a). By exploiting the vectorization properties [^2] Given any $X_{1}\in\mathbb{R}^{n_{1}\times n_{2}}$, $X_{2}\in\mathbb{R}^{n_{2}\times n_{3}}$, and $X_{3}\in\mathbb{R}^{n_{3}\times n_{4}}$, it holds $\mathtt{vec}\left(X_{1}X_{2}X_{3}\right)=(X_{3}^{\top}\otimes X_{1})\mathtt{vec}\left(X_{2}\right)$., we can manipulate (A.4) to obtain the system
$\mathtt{vec}\!\left(\mathrm{w}_{{t+1}}\mathrm{w}_{{t+1}}^{\top}\right)\!=\!(F\otimes F)\mathtt{vec}\left(\mathrm{w}_{t}\mathrm{w}_{t}^{\top}\right)$ (A.5a)
$H^{\text{vc}}_{t+1}\!=\!\lambda H^{\text{vc}}_{t}\!+\!(M\otimes M)\mathtt{vec}\left(\mathrm{w}_{t}\mathrm{w}_{t}^{\top}\right)$ (A.5b)
$S^{\text{vc}}_{t+1}\!=\!\lambda S^{\text{vc}}_{t}\!+\!({\theta^{\star}}^{\top}\!M\!\otimes\!M)\mathtt{vec}\left(\mathrm{w}_{t}\mathrm{w}_{t}^{\top}\right)\!.$ (A.5c)
We note that (A.5) describes a cascade of a linear system with state $\mathtt{vec}\left(\mathrm{w}_{t}\mathrm{w}_{t}^{\top}\right)$ feeding a linear system with state $(H^{\text{vc}}_{t},S^{\text{vc}}_{t})$. Hence, it is well known that its steady-state can be characterized by resorting to a Sylvester equation. To this end, let $\Pi_{H}\in\mathbb{R}^{(n+m)^{2}\times n_{\mathrm{w}}^{2}}$ and $\Pi_{S}\in\mathbb{R}^{(n+m)n\times n_{\mathrm{w}}^{2}}$ solve
$\Pi_{H}(F\otimes F)$ $=\lambda\Pi_{H}+M\otimes M$ (A.6a)
$\Pi_{S}(F\otimes F)$ $=\lambda\Pi_{S}+({\theta^{\star}}^{\top}M)\otimes M.$ (A.6b)
Since $F$ is marginally stable (cf. Assumption 3.1) and $\lambda\in(0,1)$, then $\sigma(F\otimes F)\cap\sigma(\lambda I)=\emptyset$. Thus, the solutions $\Pi_{S}$ and $\Pi_{H}$ to (A.6) exist and are unique (see, e.g., [56]). The proof of (27) follows by (i) noticing that $(\Pi_{\mathrm{x}},\Pi_{H},\Pi_{S})$ are used to define $\chi^{{\textnormal{ss}}}$ (cf. (26)), and (ii) plugging (A.2) and (A.6) into system (24).
As for (28), it can be shown using an algebraic manipulation of (A.6). Indeed, by pre-multiplying (A.6a) by ${\theta^{\star}}^{\top}\otimes I_{n+m}$ we can write
$({\theta^{\star}}^{\top}\!\otimes\!I_{n+m})\Pi_{H}(F\otimes F\!-\!\lambda I)$ $\!=\!({\theta^{\star}}^{\top}\otimes I_{n+m})(M\otimes M)\!\stackrel{{\scriptstyle(a)}}{{=}}\!({\theta^{\star}}^{\top}M)\otimes M$
=$\!\stackrel{{\scriptstyle(b)}}{{=}}\!\Pi_{S}(\!F\!\otimes\!F\!-\!\lambda I),$ (A.7)
where in $(a)$ we used the mixed-product property of the Kronecker operator [^3] Given any $X_{1}\!\in\!\mathbb{R}^{n_{1}\times n_{2}}$, $X_{2}\!\in\!\mathbb{R}^{n_{3}\times n_{4}}$, $X_{3}\!\in\!\mathbb{R}^{n_{2}\times n_{5}}$, and $X_{4}\!\in\!\mathbb{R}^{n_{4}\times n_{6}}$, it holds $(X_{1}\!\otimes\!X_{2})(X_{3}\!\otimes\!X_{4})\!=\!(X_{1}X_{3})\!\otimes\!(X_{2}X_{4})$., while $(b)$ follows from (A.6b). So that the proof is complete.